# Are LLMs Better than Reported? Detecting Label Errors and Mitigating Their Effect on Model Performance

### v1

@@ 4.1 Do current benchmarks include mislabeled data?
To address the first research question, we conducted a detailed analysis of current benchmarks, identifying the extent to which mislabeled data exists across various datasets. Following the procedure described in subsection 3.2, we annotate the test-set using LLMs. We then contrast these annotations with the original labels, to find disagreements. As shown in Table 2, the disagreement rate is significant and can be up to $\sim 40\%$ of the examples. For instance, for 34.4% of the examples in the BEGIN dataset, the LLM ensemble label differ from the original label. Example for such label error detection is presented in subsection 4.1. While usually we would say that this means that the LLMs performed poorly, we choose to further investigate these examples and settle the disagreement. To this end, we asked human experts to re-annotate these examples (as described in subsection 3.2), without any knowledge on LLM’s $p$ or the original label. After experts re-annotation, we can conclude which is correct: the original label, or the LLMs?
Our findings show a considerable number of label errors for all examined datasets (see the %error column in Table 2). Based on the experts Gold label and the sample sizes, we also estimate a lower bound for the total percentage of label errors in the full datasets. For this calculation, we employed the Clopper-Pearson exact method (Clopper & Pearson, 1934) to construct a 95% confidence interval for the binomial proportion, adjusted by a finite population correction (FPC) (see more details in Appendix, D.1). We provide the lower bound of these confidence intervals in parenthesis Table 2, in the %error column. For instance, we bounded the %error in MNBM to be at least 11.6%.
@@
Table 2: Summary of LLM disagreement and label error rates across different datasets. %pos is the percentage of positive (i.e., the consistent class) examples in the data. % LLM disagree refers to the percentage of examples where the LLM label differs from the original one. % error indicates the error rate in the sampled test set, while the number in parentheses denoting the estimated lower bound of the error rate for the entire dataset.
@@
4.2 LLMs for Label Error Detection
In this subsection, we explore the role of LLMs in label error detection and emphasize the strengths of an ensemble approach for this purpose.
4.2.1 Can LLMs Automatically Detect Label Errors?
@@
As described in subsection 4.1, we utilize LLMs to flag candidates for mislabeling, and indeed find label errors. In this subsection, we focus on the LLM viewpoint, including the effect of LLM confidence, and the power of ensemble (i.e., aggregating multiple LLMs). LLM annotations are valuable for flagging mislabeled data, offering more than just binary labels. By considering LLM confidence scores alongside their predictions, we can improve the precision of automatic error detection.
To better assess the utility of LLMs in detecting label errors, we break down the flagged examples into confidence-based bins. The rationale behind this is that not all flagged examples should be treated equally. Some instances are flagged with lower confidence, indicating that the LLM recognizes a potential issue but is uncertain about the label it suggests. On the other hand, when the LLM is highly confident in assigning a label that opposes the original one, it serves as a stronger signal of a possible label error.
@@
Figure 2: When LLMs disagrees with original labels - who is correct?. As the LLM’s confidence grows, so does the precision of identifying an error in the original labels.
@@
Figure 2 shows the rate of the experts agreement with the LLMs compared to the agreement with original labels, divided into confidence-based bins. Each bin includes at least 35 examples, and is defined by a confidence interval of 95% based on bootstrap sampling (see Appendix D.2 for futher details). The bins reflect increasing levels of LLM confidence in its predicted label (i.e., a stronger disagreement between LLMs and the original labels).
@@
From the figure, we observe a clear trend: as LLM confidence increases, so does its precision in detecting label errors in the original dataset. In the highest confidence bin, LLM annotations surpass the original labels in agreement with expert re-labeling, and this difference is statistically significant. This indicates that when the LLM is highly confident in its disagreement with the original label, the labeled example serves as a strong candidate for a labeling error. Note that even in cases where the expert agreement with LLMs were below 50%, mislabeled data was still discovered.
4.2.2 The power of ensemble
Figure 3: LLM Ensemble of different sizes (random subsets). (Left) presents the performance of the ensemble in terms of ROC AUC compared to the gold labels. (Right) presents the increasing ability to detect label errors. F1 is computed over Error / Not Error predictions.
As mentioned in subsection 3.2, we treat the LLM annotations as an ensemble of 2 models combined with 4 different prompts, in order to ensure greater stability in the results. Where one LLM may succeed, the other may fail, and averaging all their probabilities enables us to have more confidence in the final answer. In this subsection, we further analyzed the performance of LLMs by varying the size of the LLM ensemble, examining how this affects the model performance. We evaluate two aspects of model performance. First, we assess how closely the ensemble’s annotations match the gold labels — essentially, how much we can trust the LLM annotations. We measure this aspect of label quality using the ROC AUC compared to the gold labels. The second aspect is the ensemble’s ability to detect label errors. For this, we compute the F1-score by averaging the recall of errors and the precision of correctly identifying a candidate as a true error.
@@ 4.2.2 The power of ensemble
Results are shown in Figure 3. For both aspects, we see a clear trend. As we increase the number of models in the ensemble, the performance increases. In terms of ROC AUC w.r.t the gold labels (left plot), this suggests better annotation quality, while the right plot, a higher F1 score indicates a stronger error detector, either by recalling more errors or improving precision, or through a balance of both. Additionally, for both measures the variance decreases as the ensemble size grows, which indicates more stable and consistent annotations and error detections. Although not yet discussed in the context of error detection with LLMs, these results align with previous work showing the power of ensemble (Dietterich, 2007). These observations justify our choice to use an ensemble of models rather than a single one.
5 Comparing Annotation Approaches
Figure 4: Annotation approaches comparison.
As discussed in section 2, we focus on three main annotation approaches, each with its own benefits and drawbacks. These approaches differ in how they manage the trade-offs between label quality, scalability, and cost. In the following section, we discuss and compare their characteristics. A summary of this comparison is given in Figure 4.
5.1 Annotation Quality
Figure 6 presents the F1-score between each pair of annotation approaches. As the figure shows, LLMs have disagreements with the original labels (0.72). Yet, as discussed in subsection 4.1, the original labels themselves contain mistakes, so this disagreement does not necessarily indicate poor performance of the LLMs. When considering the Gold as the true label, LLM performance increases to 0.83. This suggests that LLMs, despite their discrepancies with the original labels, perform closer to the truth than initially reported. The Gold label, obtained by experts, has high agreement with both the Original and LLM labels. On the other hand, the MTurk-Majority approach performs poorly, with near-random F1-scores compared to both the original and gold labels, and even when compared to its stricter variant, MTurk-Strict. The results indicate that basic crowd-sourcing, without additional training to enhance crowd-workers into specialized sub-experts, performs significantly worse compared to other approaches, including LLM-based methods.
Figure 5: Comparison between all annotation methods, measured by the weighted-F1-score. Rows represent the ”true” label and columns represent the ”prediction”. For instance, the score of LLMs compared to the Original label is 0.72.
Crowd-sourcing.  For crowd-sourcing, the reported F1-score does not provide the complete picture. When we focus on individual annotators, we see that those who annotate more examples generally deliver higher-quality annotations, achieving greater accuracy when compared to both the original and gold labels (see Figure 6). This phenomenon can be explained by two hypotheses: (1) a learning process— as the annotators see more examples, they improve at the task, or (2) users who dedicate time to annotating multiple examples are likely those who either read the guidelines carefully and strive to perform the task to the best of their ability, or are naturally proficient at the task and therefore continue annotating. Even though annotators who label more instances tend to provide higher-quality annotations, they are less common—most annotators tend to stop after only a few examples. This distribution of annotators results in overall insufficient annotation quality. Pre-qualification tests are often used to shift this distribution from the ”average worker” towards more experienced or dedicated annotators; however, this requires a significantly larger budget and greater micro-management involvement from the researcher.
5.2 Consistency
Usually, when annotating a dataset, more than one annotator is involved. This applies to crowd-workers, experts, and even LLMs— in this study, we use an ensemble of different LLMs and prompts. The use of multiple annotators, similar to an ensemble, is meant to overcome the variance between individuals, which can arise from the subjective nature of NLP tasks, different interpretations of instructions, lack of experience, task difficulty, and cognitive bias (Uma et al., 2021).
As such, a common practice in the NLP community is to report Inter Annotator Agreement (IAA)— a set of statistical measures used to evaluate the agreement between individuals. Typically, IAA can be viewed as an adjustment of the proportion of pairwise agreements, where 0.0 indicates random agreement. We focus on Fleiss’s $\kappa$ (Fleiss, 1971), as it accounts for label imbalance and multiple ($>2$) annotators. High IAA, or low variance between independent annotators, is considered an indicator of high-quality annotation. In Table Table 3, we report the agreement between annotators across different approaches. For LLMs, we report two variants: (1) same model, different prompts; and (2) different models, where each model’s result is the aggregation across prompts. For reference, we also include the IAA from the original annotations, as reported in the original papers: MNBM reported an average Fleiss’s $\kappa$ of 0.696 for the hallucination annotation task; BEGIN reported Krippendorff’s $\alpha$ (a generalization of Fleiss’s $\kappa$) of 0.7; VitaminC reported Fleiss’s $\kappa$ of 0.7065 on a sample of 2,000 examples; and PAWS reported a 94.7% agreement between a single annotator’s label and the majority vote on the Wikipedia subset used in TRUE.
Table 3: Inter-Annotator Agreement in different Annotator groups. %agreement is the proportion of pairwise annotator comparison. Fleiss’s $\kappa$ (disagree. subset) refers to the $\kappa$ over the subset of disagreement between LLM and the original label.
Annotator group Fleiss’s $\kappa$agreement #examples Fleiss’s $\kappa$
@@ 5.2 Consistency
Experts 222 2    Before reconciliation 0.486 75.7 0.486    After reconciliation 0.851 93.2 0.851 MTurk 0.074 60.5 400 -0.004 3111Multiple MTurk workers have participated in annotation, yet exactly 3 annotations per example were obtained. Annotators independence assumption was made to calculate Fleiss’s $\kappa$ as with 3 annotators. LLM (different prompts) 640 4    GPT-4 0.706 85.3 0.571    PaLM2 0.750 87.7 0.696    LLaMA3 0.219 71.7 0.078    Mistral 0.459 73.2 0.314 LLMs (different models) 0.521 77.5 640 0.389 4
Experts.  While it’s true that reconciliation naturally leads to increased agreement, the significant improvement in IAA we observed highlights its importance. Though this phase is less common in practice, it is crucial not only for increasing agreement but also for improving the overall quality of annotations and ensuring more reliable outcomes. Interestingly, label changes in this phase were not symmetric, as most changes (69.3%) were in the direction of consistent $\rightarrow$ inconsistent, where one annotator found an inconsistency that the other did not (see all change details in Appendix 6). It is important to note that the $\kappa$ obtained by the experts (both before and after reconciliation) was calculated on a more challenging subset, where the original label differed from the LLM prediction, and should be interpreted with this context in mind. This is reflected in the decrease in $\kappa$ observed for all other annotator groups on this subset.
LLMs.  GPT-4 and PaLM2, the better-performing LLMs on this task, show high IAA, with $\kappa=0.706$ and $\kappa=0.75$, respectively, which is similar to the experts’ reported $\kappa$. This suggests a comparable level of variance and quality in annotation, providing further empirical evidence for considering LLMs as annotators. This property adds to previous studies showing LLMs’ quality as surrogates for human preferences (Zheng et al., 2023) or evaluations (Chiang & Lee, 2023).
Crowd-Sourcing.  Crowd workers showed near-random agreement ($\kappa=0.074$), indicating poor-quality annotations. Only 40.8% of the examples were labeled unanimously, while the rest included annotations from both classes. Even among the subset of examples unanimously labeled as consistent, 37.9% are labeled as inconsistent in both original and gold labels, pointing to a lack of attention and thoroughness. See more details on the crowd-sourced annotation distribution in Appendix A.1.2.
5.3 Cost and Scalability
@@ 6.1 Training on Mislabeled Data
Training on mislabeled data can harm model performance and stability, as learning from errors makes it harder to identify consistent patterns. The impact depends on various factors , such as the fraction of mislabeled data and the training procedure. In this subsection, we show that addressing this issue, even heuristically, significantly improves model’s performance on a test set.
Handling Label Errors.  In order to handle label errors in the training set, and reduce its effect on model performance, we propose two manipulations. For both manipulations, we use similar methodology as discussed in this paper, and flag examples where the LLMs $p$ strongly disagree (i.e., above a certain confidence threshold) with the original label. The first manipulation is filtering flagged examples out, which maintains a ”cleaner” yet smaller training set. The second manipulation is label flipping for flagged examples, which maintains the same amount of data, but may also cause harm if flipping too many correct labels.
Experimental Setup.  We set the training set to be the additional data examples from the datasets (i.e., MNBM, BEGIN, VitaminC, PAWS), which are disjoint from the test set. Note that we posses gold labels for the test set alnoe, while for the training set we only extract $p$ via GPT4 and PaLM2. The finetuning procedure includes splitting the training set into train and validation sets, and fine-tuning on the train set. We report results averaged over five seeds.
As an ablation study, we also apply these manipulations on a random subset of examples rather than the flagged examples. TThe ablation study aims to maintain a consistent number of training examples, while the ablation for flipping aims to address the claim that in some cases, a relatively small fraction of label errors may be even considered as a noise that improves model robustness (e.g., as in label perturbation (Zhang et al., 2018) or label smoothing (Szegedy et al., 2016)).
We conducted this experiment starting from two base models: DeBERTa-base-v3, [^3] https://huggingface.co/microsoft/deberta-v3-base and a fine-tuned version of it on classic NLI datasets, which we will refer to as the NLI-base model [^4] https://huggingface.co/MoritzLaurer/DeBERTa-v3-base-mnli-fever-anli. We chose the NLI-base model as NLI tasks closely resemble factual consistency evaluation (FCE), making it well-suited for this experiment. Given the similar trends, we present the results for the NLI model here. Additional experiments and implementation details can be found in Appendix C.1.
@@
Results.Figure 7 shows the results of our experiments. In our confidence-based approaches, we clearly see the trend that as the confidence threshold—according to which our manipulations are applied—grows, our manipulation results in improved ROC AUC for both models. This trend eventually (i.e., for high enough LLM confidence) brings these approaches to significantly outperform the baseline. In contrast, when we applied our manipulations on random subsets, we generally see a diminishing effect of manipulation, converging to the no-manipulation baseline.
@@ 6.2 Evaluating on Mislabeled Data
Experimental Setup.  To test this assumption, we evaluate the performance of nine models, mostly state-of-the-art LLMs, on the test datasets. We compare their performance between the Original labels, and the Gold labels. For LLMs, we used zero-shot prediction as described in subsection 2.2, and averaged over prompts. For DeBERTa-based models we used the fine-tuned models from subsection 6.1, and averaged over seeds.
Results.  Prior to this work, an evaluation of these models would induce the values and ranking as in Table 4 under the Original sub-columns. However, as shown before, these datasets include labeling errors, and therefore do not support fair evaluation. Considering the new gold labels, based on expert intervention (as described in subsection 3.2), we obtain different results, shown in the Gold sub-columns. The first observed discrepancy is the ranking of models. For example, DeBERTa-v3 has shifted from being the second-best to the second-worst. Beyond the change in ranking, all metrics’ (i.e., ROC AUC, F1-score and accuracy) range has shifted upward, indicating that LLMs perform better on this task than what was previously thought, likely due to label errors. If this phenomenon extends to other tasks and datasets beyond those examined in this study, it could suggest that LLMs are better than currently perceived.
7 Related Work
7.1 LLMs in the Annotation Loop
LLMs have been increasingly utilized as annotators in various NLP tasks, offering potential benefits in efficiency and scalability. Several studies have demonstrated that LLMs can effectively generate annotations from scratch, sometimes outperforming human annotators or crowd workers (He et al., 2023; Gilardi et al., 2023; Törnberg, 2023; Calderon & Reichart, 2024). However, LLMs are not flawless and cannot be considered gold-standard annotators when used alone. They may produce incorrect annotations, especially in complex (Chen et al., 2024), social (Ventura et al., 2023; Felkner et al., 2024), emotional (Lissak et al., 2024), or low-resource (Bhat & Varma, 2023) contexts. These studies showed that LLMs can exhibit poor performance and biases, highlighting the necessity of human oversight to ensure quality or fairness. To address this issue, several approaches for collaborative (Kim et al., 2024; Li et al., 2023) or active learning (Zhang et al., 2023; Kholodna et al., 2024) were suggested, where LLMs and humans are both part of the annotation procedure.
7.2 Handling Label Errors
Label errors (also referred to as label noise) in training and evaluation datasets can significantly impair NLP model performance and reliability (Frénay & Verleysen, 2014). Fine-tuned models have been employed to detect mislabeled data by identifying examples with high loss or low confidence (Chong et al., 2022a; Hao et al., 2020; Pleiss et al., 2020; Northcutt et al., 2019). For example, Chong et al. (2022a) demonstrated that fine-tuned pre-trained language models can effectively detect label errors by ranking data points based on the training loss. Once these high-loss or low-confidence examples are flagged, they are typically filtered out (Nguyen et al., 2019; Northcutt et al., 2019), corrected automatically (Pleiss et al., 2020; Hao et al., 2020), or re-labeled by human annotators (Northcutt et al., 2021) to verify and improve dataset quality.
7.3 Factual Consistency Evaluation
Factual consistency evaluation (FCE) refers to the task of verifying that generated text remains true to the facts in the source content, addressing factual inaccuracies in models’ outputs. It has been applied to tasks like summarization (Kryscinski et al., 2019; Xie et al., 2021; Gekhman et al., 2023) and dialogue (Honovich et al., 2021; Xue et al., 2023), which are prone to suffer from hallucinated outputs. Benchmarks like the TRUE Honovich et al. (2022) standardize evaluation across datasets. Common methods include entailment-based models (Laban et al., 2022) and QA-based approaches such as Q2 (Honovich et al., 2021) and QAFactEval (Fabbri et al., 2021), with recent advancements like WeCheck (Wu et al., 2023) improving efficiency through weakly supervised learning.

### v2

@@ 5.1 Do current benchmarks include mislabeled data?
To address the first research question, we annotate the test-set of TRUE (as described in section 4 using LLMs. We then contrast these annotations with the original labels, to find disagreements. As shown in Table 2, the disagreement rate is significant and can be up to $\sim 40\%$ of the examples. An example of such disagreement is presented in Table 1. While this would typically suggest poor LLM performance, we further investigated by re-annotating with experts to determine which was more accurate: the original label or the LLMs’ prediction.
Our findings show a considerable number of label errors for all examined datasets (see the %error column in Table 2). Based on the experts gold label and the sample sizes, we also estimate a lower bound for the total percentage of label errors in the full datasets. We employed the Clopper-Pearson exact method (Clopper and Pearson, 1934) to construct a 95% confidence interval for the binomial proportion, adjusted by a finite population correction (FPC) (see more details in Appendix G.1). We provide the lower bound of these confidence intervals in parentheses in Table 2, under the %error column. The lower bounds range from 3% in the PAWS dataset to 15.8% in the BEGIN dataset.
@@
Table 2: Summary of LLM disagreement and label error rates across different datasets. %pos is the percentage of positive (i.e., the consistent class) examples in the data. % LLM disagree refers to the percentage of examples where the LLM label differs from the original one. % error indicates the error rate in the sampled test set, while the number in parentheses denotes the estimated lower bound of the error rate for the entire dataset.
5.2 Can LLMs Detect Label Errors?
@@
As described in subsection 5.1, we utilize LLMs to flag candidates for mislabeling, and indeed find label errors. In this subsection, we focus on the LLM viewpoint, exploring the effect of LLM confidence, and the power of ensemble.
Confidence LLM annotations are valuable for flagging mislabeled data, offering more than just hard labels. By considering LLM confidence scores alongside their predictions, we can improve the precision of automatic error detection. Leveraging confidence can reduce re-annotation efforts by flagging only cases exceeding a predefined threshold. The rationale is that not all flagged examples should be treated equally. Instances flagged with low confidence indicate that the LLM recognizes a potential issue, however, when the LLM is highly confident in a label that contradicts the original one, it provides a stronger signal of a possible error.
@@
Figure 2: When LLMs disagree with original labels - who is correct? (Top) TRUE (Bottom) SummEval. As the LLM’s confidence grows, so does the precision of identifying an error in the original labels.
@@
Figure 2 shows the rate of the experts’ agreement with the LLMs compared to the agreement with original labels, divided into confidence-based bins. Bins are balanced by size, and defined by a confidence interval of 95% based on bootstrap sampling (see Appendix G.2 for further details). The bins reflect increasing levels of LLM confidence in its predicted label (i.e., a stronger disagreement between LLMs and the original labels).
@@
From the top of Figure 2, we observe a clear trend: as LLM confidence increases, so does its precision in detecting label errors in the original dataset. In the highest confidence bin, LLM annotations surpass the original labels in agreement with expert re-labeling, and this difference is statistically significant. This indicates that when the LLM is highly confident in its disagreement with the original label, the labeled example serves as a strong candidate for a labeling error. Note that even in cases where the expert agreement with LLMs was below 50%, mislabeled data was still discovered. See Appendix E for model-specific analysis.
We replicated this analysis on the SummEval dataset (bottom of Figure 2) and observed a similar trend: higher confidence increases the likelihood that the LLM prediction is closer to the expert annotation than the original label. In the SummEval case, we consider the crowd-sourced labels as the original labels. For more details see Appendix A.
Ensemble By varying the size of the LLM ensemble, we examine two key aspects: predictive power (how well predictions align with gold labels, measured by ROC AUC for TRUE and average correlation for SummEval), and error detection power (measured by F1-score, averaging the recall of errors and the precision of correctly identifying a candidate as a true error). The ensemble power analysis is presented in Figure 3.
Figure 3: The power of ensemble. (Top) TRUE (Bottom) SummEval. As the ensemble size increases (x-axis), its performance against gold labels (Left), and its ability to detect label errors (Right) improve.
@@
For both aspects, we see a clear trend. As we increase the number of models in the ensemble, the performance increases. A higher ROC AUC with respect to the gold labels (left) reflects better annotation quality, while a higher F1 score (right) indicates a stronger error detector, either by recalling more errors or improving precision, or through a balance of both. Notably, to place the absolute F1-score in context, the expected F1-score for random behavior is approximately 0.22 (when randomly flagging errors), or around 0.13 (when randomly guessing the annotation), due to the class imbalance between error and non-error cases. Additionally, for both measures, the variance decreases as the ensemble size grows, which indicates more stable and consistent annotations and error detections. Similarly, Figure 3 (bottom) shows the power of LLM ensemble on the same aspects on the SummEval datasets, aggregated over four summarization dimensions (see experiment details on Appendix A.4.2). Trends of diminishing variance and increased performance and error detection are observed here as well.
Although not yet discussed in the context of error detection with LLMs, these results align with previous work showing the power of ensemble (Dietterich, 2007).
Our findings show that incorporating multiple LLMs and prompts in an ensemble is valuable: as the ensemble size increases, both label quality and error detection improve. These observations justify our choice to use an ensemble of models rather than a single one.
Annotator group Fleiss’s $\kappa$ %agreement #examples Fleiss’s $\kappa$
@@
Experts 222 2    Before reconciliation 0.486 75.7 0.486    After reconciliation 0.851 93.2 0.851 MTurk 0.074 60.5 400 -0.004 3111Multiple MTurk workers have participated in annotation, yet exactly 3 annotations per example were obtained. Annotator independence assumption was made to calculate Fleiss’s $\kappa$ as with 3 annotators. LLM (different prompts) 640 4    GPT-4 0.706 85.3 0.571    PaLM2 0.750 87.7 0.696    LLaMA3 0.219 71.7 0.078    Mistral 0.459 73.2 0.314 LLMs (different models) 0.521 77.5 640 0.389 4
Table 3: Inter-annotator agreement (IAA) across annotator groups. LLMs such as GPT-4 and PaLM2 approach expert-level agreement, while MTurk workers show low and inconsistent reliability. Results for SummEval are provided in Table 5 in the appendix.
6 Comparing Annotation Approaches
Our paper discusses three annotation approaches, each with its own benefits and drawbacks, differing in how they balance label quality, scalability, and cost. Here we summarize the main findings, with additional analyses provided in Appendix B. Figure 4 highlights the key results.
Figure 4: Annotation approaches comparison.
LLMs exhibit strong agreement with experts and among themselves. Inter-annotator agreement (IAA) among LLMs, as well as their alignment with expert annotations, are significantly higher than that of crowd workers. As shown in Table 3, GPT-4 and PaLM2 achieve $\kappa$ scores above 0.70, approaching expert-level agreement after reconciliation ($\kappa=0.85$). In contrast, MTurk workers reach only $\kappa=0.07$, underscoring the gap between crowd- and LLM-based annotation.
Crowd worker quality improves with experience but remains inconsistent. Our analysis shows that experienced crowd workers produce higher-quality annotations, as illustrated in Figure 5. However, even among them, annotation quality and consistency remain lower than LLM-based annotation, which is more reliable. This is reflected in the wide variance of MTurk agreement (60.5% overall, $\kappa=-0.004$ on disagreement cases), suggesting that crowd annotation requires substantial verification to ensure reliability.
LLMs provide fast, scalable, and cost-efficient annotation. Compared to expert and crowd-sourced annotation, LLMs require less time and are much more cost-effective per annotation. As discussed in subsection B.3, LLM annotation is estimated to be 100-1000 times cheaper than human annotation. This makes them a viable alternative for large-scale annotation while effectively balancing the trade-off.
@@ 7.1 Training on Mislabeled Data
Training on mislabeled data can harm model performance and stability, as learning from errors makes it harder to identify consistent patterns. The impact depends on various factors, such as the fraction of mislabeled data and the training procedure. In this subsection, we show that addressing this issue, even heuristically, significantly improves the model’s performance on a test set.
Handling Label Errors In order to handle label errors in the training set, and reduce its effect on model performance, we propose two manipulations. For both manipulations, we flag examples where the model strongly disagrees with the original label(i.e., with confidence above a certain threshold). The first manipulation is filtering flagged examples out, which maintains a “cleaner” yet smaller training set. The second manipulation is label flipping for flagged examples, which maintains the same amount of data, but may also cause harm if flipping too many correct labels.
@@
Experimental Setup We set the training set to be the additional data examples from the datasets (i.e., MNBM, BEGIN, VitaminC, PAWS), which are disjoint from the test set. Note that we posses gold labels for the test set alone, while for the training set we only extract the confidence. The finetuning procedure includes splitting the training set into train and validation sets, and fine-tuning on the train set. We report average results of five seeds.
As an ablation study, we also apply these manipulations on a random subset of examples rather than the flagged examples. The ablation study aims to maintain a consistent number of training examples, while the ablation for flipping aims to address the claim that in some cases, a relatively small fraction of label errors may be even considered as a noise that improves model robustness (e.g., as in label perturbation (Zhang et al., 2018) or label smoothing (Szegedy et al., 2016)).
We conducted this experiment starting from two base models: DeBERTa-v3, and a fine-tuned version of it on classic NLI datasets, which we will refer to as the NLI-base model. We chose the NLI-base model as NLI tasks closely resemble factual consistency evaluation (FCE), making it well-suited for this experiment. Given the similar trends, we present the results for the NLI model here. Additional experiments and implementation details can be found in Appendix F.1.
Results Figure 6 shows the results of our experiments. In our confidence-based approaches, we clearly see the trend that as the confidence threshold, according to which our manipulations are applied, grows, our manipulation results in improved ROC AUC for both models. This trend eventually (i.e., for high enough LLM confidence) brings these approaches to significantly outperform the baseline. In contrast, when we applied our manipulations on random subsets, we generally see a diminishing effect of manipulation, converging to the no-manipulation baseline.
@@ 7.2 Evaluating on Mislabeled Data
Experimental Setup To test this assumption, we evaluate the performance of nine models, mostly state-of-the-art LLMs, on the test datasets. We compare their performance between the original labels, and the gold labels. For LLMs, we used zero-shot prediction as described in section 3, and averaged over prompts. For DeBERTa-based models, we used the fine-tuned models from subsection 7.1, and averaged over seeds.
Results Prior to this work, an evaluation of these models would induce the values and ranking as in Table 4 under the Original sub-columns. However, as shown before, these datasets include labeling errors, and therefore do not support fair evaluation. Considering the new gold labels, based on expert intervention (as described in subsection 4.2), we obtain different results, shown in the Gold sub-columns. The first observed discrepancy is the ranking of models. For example, DeBERTa-v3 has shifted from being the second-best to the second-worst. Beyond the change in ranking, all metrics’ (i.e., ROC AUC, F1-score, and accuracy) range has shifted upward, indicating that LLMs perform better on this task than previously thought. We further discuss the performance differences between LLMs and fine-tuned models in Appendix F.2. If this phenomenon extends to other tasks and datasets beyond those examined in this study, it could suggest that LLMs are better than currently perceived.